Math:50:640:042 Intermediate Algebra Spring 2010

Course Specifics:

Instructor: Tejas Gandhi Office: 413 Business and Science Building

E-mail:  tejasg@camden.rutgers.edu

Classroom: ATG 101 Office Hours: Tue 5:00-5:45pm

Fri 10:05-10:50am

If you have any questions and cannot make it to my office hours please send me email and we can arrange to meet.

Text Book:

The required text for this course is the following. Elementary and Intermediate Algebra, 5^{th} edition, Jerome Kaufman.

Course Objectives:

The basic requirement upon completion of this couses, is the student’s ability to apply fundamental math concepts. In performing this task, the student will demonstrate a proficiency in solving equations (one variable) and inequalities, operations with polynominals, rational expressions involving exponents (integers and rational), solving system of linear equations, graphing and generating equations of lines, solving systems of equations of lines, solving quadratic equations and graphing parabolas.

General Syllabus:

The following chapters will be covered: 1, 2, 3, 4, - Reviews 5, 6, 7, 9.1-9.5, 10, 11.2 and 11.4, 12.2-12.4

Grading:

Your final grade in the course will be decided by your performance in 4 exams.

1. Exam 1 – 25%.

50:640:042 2

2. Exam 2 – 25%.

3. Exam 3 – 25%.

4. Comprehensive final – 25%. The final exam is mandatory (required by the State). Missed exams cannot be made up unless arrangements are made prior to the exam date.

Attendance:

There are no excused absences. Attendance of all classes is expected. Any material or announcements missed is the full responsibility of the student.